find the midpoint of the segment with the endpoints (square root of 96 and square root of 3) and ( square root of 6 and the square root of 12)
Answer by checkley75(3666)

(sqrt96-sqrt3)/2=(8-1.732)/2=6.2679/2=3.134 is the mid point.
-------------------------------------------------------------------
(sqrt12-sqrt6)/2=(3.464-2.45)/2=1.014/2=.507